OpenStudy (jmark):

Without using trigonometric tables,find the value of cos 70/sin 20+cos 57.cosec 33-2 cos 60

OpenStudy (kc_kennylau):

(I have no idea at all) try substituting cos 70 with sin(90-20) and cos57 with sin(90-33)?

OpenStudy (raden):

cos 70/sin 20 + cos 57.cosec 33 - 2 cos 60 = sin20/sin20 + sin33 * 1/sin33 - 2 cos 60 = 1 + 1 - 2 * 1/2 = 1 + 1 - 1 = 1